OpenStudy (anonymous):

can anyone help me wit algebra. (-3a^2+6a+9)-(6a^2+5a+7)

OpenStudy (anonymous):

\[-9a^2+a+2=0\] use quadratic formula \[a= \frac{ -b \pm \sqrt{(b)^2-4ac} }{ 2a }\] \[a1=\frac{ 1-\sqrt{73} }{ 18 }\]\[a2=\frac{ 1+\sqrt{73} }{ 18 }\]

OpenStudy (anonymous):

Multiply the negative to (6a^2+5a+7) so that it equals \[-6a ^{2}-5a-7\] Then , you subtract: \[-3a ^{2}+6a+9-6a ^{2}-5a-7\] you get \[-9a ^{2}+a+2\]